The two baskets of fruits A and B weigh 100 kilograms. Now take out 12 kilograms of fruits from tube A and put them in basket B. The fruits in basket A are 4 kg heavier than basket B. How many kilograms do the two baskets of fruit weigh?


1
Expert's answer
2021-08-03T14:11:19-0400

Let "x=" the weight of fruits in basket A, "y=" the weight of fruits in basket B.


"x+y=100""x-12=y+12+4"

"x=y+28""y+28+y=100"

"x=64""y=36"

The basket A weighs 64 kilograms.

The basket B weighs 36 kilograms.